OpenStudy (izzy18970):

Please Help. Solve the following system. 2x+y+2z=1 8x-4y+2z=-6 2y+2z=2

OpenStudy (faiqraees):

y+2z = 1-2x -2y+z = -3-4x Eliminate y and then solve the simultaneous equation

OpenStudy (izzy18970):

What are the steps to solving the equation @FaiqRaees

OpenStudy (faiqraees):

Multiply my first equation by 2 and then add the equations

OpenStudy (sbuck98):

Izzy, do you have a smart phone? If so, download photomath, it shows you all the steps to equation problems

OpenStudy (izzy18970):

thank yous

OpenStudy (sbuck98):

Your welcome, message me when you have free time!

OpenStudy (izzy18970):

ok